OpenStudy (anonymous):

Find the perimeter of a rectangular object which has a length of sqr(75) feet and a width ofsprt(48) feet. Part 2: Explain, in complete sentences, how you arrived at the answer and give the final solution in simplified radical form. (2 points) Part 3: What type of object in your home or school might this be? (1 point)

OpenStudy (anonymous):

help me out plz

OpenStudy (anonymous):

Is the length and width \[\sqrt{75}and \sqrt{48}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

First simplify the radicals\[\sqrt{75}=\sqrt{25*3}=5\sqrt{3};\sqrt{48}=\sqrt{16*3}=4\sqrt{3}\]Next, the perimeter is 2L+2W, or\[10\sqrt{3}+8\sqrt{3}=18\sqrt{3}\]This is the perimeter. (Part 2 to follow)

OpenStudy (anonymous):

The perimeter is twice the length plus twice the width for a rectangle. The answer already is in simplified radical form. (Part 3 to follow)

OpenStudy (anonymous):

Based on the dimensions of about 8.6 and 6.9 feet, this might be a table, for example.

OpenStudy (anonymous):

thanks